INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF PLATINUM NITRIDE IN THE NEAR IN-FRARED

\underline{LEAH C. O’BRIEN}, RACHEL A. HARRIS, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; SEAN WHITTEMORE, JAMES J. O’BRIEN, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499.

A new electronic transition of PtN has been recorded using intracavity laser absorption spectroscopy. Four red-degraded branches are observed, with a bandheads located at 11733 and 11725 cm^{−1}. The results of the analysis will be presented and compared with ab initio calculations.